you will mix 80% acetone with 30% acetone to make 40 ml of 50% acetone.how much of each starting solution do you need?
Answer by clearblueskai(22)

let x = mL of 80% acetone
40-x = mL of 30% acetone
(0.8)x+(40-x)0.3 = (40)0.5
0.8x+12-0.3x= 20
0.5x = 8
x = 16
40-x = 40-16 =24